A mixture of Helium and Neon gases is collected over water at 28 degree Celsius and 745 mm Hg. If the partial pressure of Helium is 368 mm Hg, what is the partial pressure of Neon? Please explain.
1
Expert's answer
2015-07-13T05:51:04-0400
The vapor pressure at 28 Co is of 28.3 mm Hg. Based on the Dalton’s law regarding partial pressure of gas mixture the total pressure is defined as follows:
P = p(Ne) + p(He) + p(H2O),
thus, partial pressure of Neon equals:
p(Neon) = P - p(He) - p(H2O) = 745 mm Hg – 368 mm Hg – 28.3 mm Hg = 348.7 mm Hg